# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ 2.2 Asymptotic properties
    $\left(\bm{\dot{\mu}(\beta_{0})}^{T}\bm{\dot{\mu}(\beta_{0})}\right)^{1/2}\left(\bm{\widehat{\beta}_{\alpha}}-\bm{\beta_{0}}\right)\xrightarrow{\mathcal{D}}\mathcal{N}_{p}\left(\bm{0}_{p},\ v_{1\alpha}\sigma_{0}^{2}\mathbb{I}_{p}\right),$
    and
    $\sqrt{n}\left(\widehat{\sigma}_{\alpha}^{2}-\sigma_{0}^{2}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\ v_{2\alpha}\sigma_{0}^{4}\right),$
where
$v_{1\alpha}=\frac{1}{\sigma^{2}}\frac{\zeta_{2\alpha}}{\zeta_{\alpha}^{2}}=\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{3/2},$
$v_{2\alpha}=\frac{1}{\sigma^{4}}\frac{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}{\varsigma_{\alpha}^{2}}=\frac{4}{\left(2+\alpha^{2}\right)^{2}}\left[2\left(1+2\alpha^{2}\right)\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{5/2}-\alpha^{2}(1+\alpha)^{2}\right].$
@@
Note that, these AREs are independent of the choice of the (non-linear) regression function $\mu(\cdot,\cdot)$ and exactly the same as those obtained for the linear regression model in Ghosh and Basu [14]. The exact values of the ARE of the MDPDE of $\bm{\beta}$ for different $\alpha$ are reported in Table 1, along with the same for the existing robust competitors (as described in Appendix A.2). From the table, it is evident that for the MDPDE, the loss in efficiency is very small for smaller values of $\alpha$, in comparisons to other competing estimators at suitable tuning parameter values expecting to produce similar robustness properties.
Huber’s M-est Tukey’s M-est KPS est MDPDE
c ARE c ARE w ARE $\alpha$ ARE
1.000 0.466 3.000 0.776 0.5 0.911 0.1 0.988
1.345 0.589 3.500 0.858 1.0 0.647 0.3 0.921
1.700 0.710 4.000 0.910 1.3 0.503 0.5 0.838
2.000 0.802 4.685 0.950 1.5 0.424 0.7 0.757
2.500 0.916 5.500 0.973 1.7 0.357 1.0 0.650
2.3 Influence function and sensitivity analysis
@@
Influence function (IF) is a measure of local robustness property of an estimator. Here, we will follow the general theory from Ghosh and Basu [14], as presented briefly in Appendix A.1.
@@
Firstly, the minimum DPD functional $\bm{T_{\alpha}}=\left(\bm{T_{\alpha}^{\beta}},T_{\alpha}^{\sigma^{2}}\right)$ of the parameter $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ under the setup of the NLR model (3), with true distribution $g_{i}$ and $f_{i}(\cdot;\bm{\theta})\in\mathcal{F}_{i}$, is defined by
[…]
@@
where $\bm{G}=(G_{1},\cdots,G_{n})^{T}$ with $G_{i}$ being the distribution function for the density $g_{i}$ for each $i=1,\ldots,n$. Then, assuming the true distributions belonging to the corresponding model families and using the general expression of IF under INH setup (as given in Eq. (38) in Appendix A.1), the IF of $\bm{T}_{\alpha}$ with contamination only in the $i_{0}$-th observation at the contamination point $t_{i_{0}}$ is obtained as
@@
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha},\bm{F}_{\bm{\theta}})=\begin{pmatrix}\frac{1}{\zeta_{\alpha}}f_{i_{0}}(t_{i_{0}};\bm{\theta})^{\alpha}\frac{(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))}{\sigma^{2}}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}}},\bm{\beta})\\ \\ \frac{1}{n\varsigma_{\alpha}}\left[\left\{\frac{(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}}{2\sigma^{4}}-\frac{1}{2\sigma^{2}}\right\}f_{i_{0}}(t_{i_{0}};\bm{\theta})^{\alpha}+\frac{\alpha}{2}\zeta_{\alpha}\right]\end{pmatrix},$
@@
where $\bm{F}_{\bm{\theta}}=\left(F_{1,\bm{\theta}},F_{2,\bm{\theta}},\ldots,F_{n,\bm{\theta}}\right)^{T}$ with $F_{i,\bm{\theta}}$ being the distribution function for the model density $f_{i}(\cdot;\bm{\theta})$. On simplification, the IFs of $\bm{T}_{\alpha}^{\bm{\beta}}$ and $T_{\alpha}^{\sigma^{2}}$ with contamination only in the $i_{0}$-th observations at the contamination point $t_{i_{0}}$, respectively, become
@@
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=\\ (1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}}},\bm{\beta}),$ (17)
and
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\sigma^{2}},\bm{F}_{\bm{\theta}})=\\ \frac{2(1+\alpha)^{\frac{5}{2}}}{n(2+\alpha^{2})}\left\{(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}-\sigma^{2}\right\}e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(\bm{x_{i_{0}}},\bm{\beta}))^{2}}+\frac{2\alpha(1+\alpha)^{2}}{n(2+\alpha^{2})}.$ (18)
@@
Further, according to the general formula given in Eq. (39) in Appendix A.1, the IFs of $\bm{T^{\beta}}_{\alpha}$ and $T_{\alpha}^{\sigma^{2}}$ with contamination in all the observations at the contamination points in $\bm{t}=(t_{1},\cdots,t_{n})^{T}$, are, respectively, given by
@@
$IF(\bm{t},\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=\\ (1+\alpha)^{\frac{3}{2}}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\sum_{i=1}^{n}\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}},$ (19)
$IF(\bm{t},T_{\alpha}^{\sigma^{2}},\bm{F}_{\bm{\theta}})=\\ \frac{2(1+\alpha)^{\frac{5}{2}}}{n(2+\alpha^{2})}\sum_{i=1}^{n}\left\{(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}-\sigma^{2}\right\}e^{-\frac{\alpha}{2\sigma^{2}}(t_{i}-\mu(\bm{x_{i}},\bm{\beta}))^{2}}+\frac{2\alpha(1+\alpha)^{2}}{(2+\alpha^{2})}.$ (20)
@@
It is easy to observe that these IFs are also bounded in each argument $t_{i},\mbox{ for }i=1,2,\ldots,n,\mbox{ for all }\alpha>0$, and unbounded at $\alpha=0$, as before.
@@
Let us now consider some IF based summary measures of robustness for our NLR model. The gross error sensitivity of the minimum DPD functional $\bm{T^{\beta}}_{\alpha}$ of $\bm{\beta}$ at the vector of model distributions $\bm{F_{\theta}}$ with contamination only in the $i_{0}$-th observation is given by
@@
$\gamma_{i_{0}}^{u}(\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=\sup_{t}\left\{\lVert IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})\rVert\right\}\\ =\left\{\begin{matrix}\frac{(1+\alpha)^{\frac{3}{2}}}{\sqrt{\alpha}}\sigma e^{-\frac{1}{2}}\lVert\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}}},\bm{\beta})\rVert,&\mbox{if}&\alpha>0,\\ \infty,&\mbox{if}&\alpha=0.\end{matrix}\right.$ (21)
But this measure does not satisfy the scale invariance property with respect to scale transformation of the components of $\bm{\beta}$. To overcome this problem, we consider a modified measure, namely the self standardized sensitivity. For contamination only in the $i_{0}$-th observation, it is given by
@@
$\gamma_{i_{0}}^{s}(\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})=\sup_{t}\left\{IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})^{T}(\bm{\Psi}_{pn}^{-1}\bm{\Omega}_{pn}\bm{\Psi}_{pn}^{-1})^{-1}IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})\right\}^{1/2}\\ =\left\{\begin{matrix}\frac{(1+2\alpha)^{\frac{3}{4}}}{\sqrt{n\alpha}}e^{-\frac{1}{2}}\left\{\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}},\beta})^{T}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i_{0}},\beta})\right\}^{1/2},&\mbox{if}&\alpha>0,\\ \infty,&\mbox{if}&\alpha=0,\end{matrix}\right.$ (22)
@@
where $\bm{\Psi}_{pn},\bm{\Omega}_{pn}$ are the $p\times p$ principal diagonal blocks of the matrices $\bm{\Psi}_{n},\bm{\Omega}_{n}$, respectively. Note that, both the above sensitivity measures are decreasing function in $\alpha\ (>0)$. So, it is evident that, with the presence of outliers, the robustness of the MDPDE of $\bm{\beta}$ increases with increasing values of $\alpha$. However, in Section 2.2 we have observed that the ARE of MDPDE of $\bm{\beta}$ (under pure data) decreases with increasing values of $\alpha$. Together, the tuning parameter $\alpha$ acts as a trade-off between efficiency and robustness of the MDPDE under NLR models, in consistence with the existing literature of DPD based procedures.
@@
For contamination in all or some observations, the gross error sensitivity and the self standardized sensitivity can similarly be defined as in (21) and (22), but replacing $IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})$ by $IF(\bm{t},\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})$ and taking supremum over all possible $t_{1}\ldots,t_{n}$. The resulting implications are again the same, and hence these details are omitted for brevity.

### v2

@@ 2.2 Asymptotic properties
    $\left(\bm{\dot{\mu}}(\bm{\beta}_{0})^{T}\bm{\dot{\mu}}(\bm{\beta}_{0})\right)^{1/2}\left(\bm{\widehat{\beta}}_{\alpha}-\bm{\beta}_{0}\right)\xrightarrow{\mathcal{D}}\mathcal{N}_{p}\left(\bm{0}_{p},\ v_{1\alpha}\sigma_{0}^{2}\mathbb{I}_{p}\right),\mbox{ and }\sqrt{n}\left(\widehat{\sigma}_{\alpha}^{2}-\sigma_{0}^{2}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\ v_{2\alpha}\sigma_{0}^{4}\right),$
    $\mbox{where }~~~~~~~~~~v_{1\alpha}=\frac{1}{\sigma^{2}}\frac{\zeta_{2\alpha}}{\zeta_{\alpha}^{2}}=\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{3/2},$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
    $\mbox{and }~~~~~~~~v_{2\alpha}=\frac{1}{\sigma^{4}}\frac{\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}}{\varsigma_{\alpha}^{2}}=\frac{4}{\left(2+\alpha^{2}\right)^{2}}\left[2\left(1+2\alpha^{2}\right)\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{5/2}-\alpha^{2}(1+\alpha)^{2}\right].$
@@
These AREs are independent of the choice of the (non-linear) regression function $\mu(\cdot,\cdot)$ and exactly the same as those obtained for the linear regression model in Ghosh and Basu [13]. The exact values of the ARE of the MDPDE of $\bm{\beta}$ for different $\alpha$ are reported in Table C.1 of Web Appendix C, along with the same for the existing robust competitors. From the table, it is evident that the loss in efficiency of the MDPDE is very small for smaller values of $\alpha$, compared to other competing estimators at suitable tuning parameter values, expecting to produce similar robustness properties.
2.3 Influence function and sensitivity analysis of the proposed MDPDE
@@
Influence function (IF) is a measure of the local robustness property of an estimator. Here, we will follow the general theory from Ghosh and Basu [13], as presented briefly in Web Appendix A.1.
@@
Firstly, the minimum DPD functional $\bm{T}_{\alpha}=\left(\bm{T}_{\alpha}^{\bm{\beta}},T_{\alpha}^{\sigma^{2}}\right)$ of the parameter $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ under the setup of the NLR model (3), with true distribution $g_{i}$ and $f_{i}(\cdot;\bm{\theta})\in\mathcal{F}_{i}$, is defined by
$\bm{T}_{\alpha}(\bm{G})=\arg\min_{\bm{\theta\in\bm{\Theta}}}\frac{1}{n}\sum_{i=1}^{n}\left[\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{\alpha}\sqrt{1+\alpha}}-\frac{1+\alpha}{\alpha}\int f_{i}^{\alpha}(y;\bm{\theta})g_{i}(y)dy\right],$ (15)
@@
where $\bm{G}=(G_{1},\cdots,G_{n})^{T}$ with $G_{i}$ being the distribution function for the density $g_{i}$ for each $i=1,\ldots,n$. Then, assuming the true distributions belonging to the corresponding model families and using the general expression of IF under INH setup (as given in Eq. (3) in Web Appendix A.1), the IF of $\bm{T}_{\alpha}$ with contamination only in the $i_{0}$-th observation at the contamination point $t_{i_{0}}$ is obtained as
@@
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha},\bm{F}_{\bm{\theta}})=\begin{pmatrix}\frac{1}{\zeta_{\alpha}}f_{i_{0}}^{\alpha}(t_{i_{0}};\bm{\theta})\frac{(t_{i_{0}}-\mu(\bm{x}_{i_{0}},\bm{\beta}))}{\sigma^{2}}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x}_{i_{0}},\bm{\beta})\\ \\ \frac{1}{n\varsigma_{\alpha}}\left[\left\{\frac{(t_{i_{0}}-\mu(\bm{x}_{i_{0}},\bm{\beta}))^{2}}{2\sigma^{4}}-\frac{1}{2\sigma^{2}}\right\}f_{i_{0}}^{\alpha}(t_{i_{0}};\bm{\theta})+\frac{\alpha}{2}\zeta_{\alpha}\right]\end{pmatrix}$,
@@
where $\bm{F}_{\bm{\theta}}=\left(F_{1,\bm{\theta}},F_{2,\bm{\theta}},\ldots,F_{n,\bm{\theta}}\right)^{T}$ with $F_{i,\bm{\theta}}$ being the distribution function of the model density $f_{i}(\cdot;\bm{\theta})$. On simplification, the IFs of $\bm{T}_{\alpha}^{\bm{\beta}}$ and $T_{\alpha}^{\sigma^{2}}$ with contamination only in the $i_{0}$-th observation at the contamination point $t_{i_{0}}$, respectively, become
@@
$IF_{i_{0}}(t_{i_{0}},\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(\bm{x}_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(\bm{x}_{i_{0}},\bm{\beta}))^{2}}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x}_{i_{0}},\bm{\beta}),$ (16)
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\sigma^{2}},\bm{F}_{\bm{\theta}})=\frac{2(1+\alpha)^{\frac{5}{2}}}{n(2+\alpha^{2})}\left\{(t_{i_{0}}-\mu(\bm{x}_{i_{0}},\bm{\beta}))^{2}-\sigma^{2}\right\}e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(\bm{x}_{i_{0}},\bm{\beta}))^{2}}+\frac{2\alpha(1+\alpha)^{2}}{n(2+\alpha^{2})}.$ (17)
@@
Further, according to the general formula given in Eq. (4) of Web Appendix A.1, the IFs of $\bm{T^{\beta}}_{\alpha}$ and $T_{\alpha}^{\sigma^{2}}$ with contamination in all the observations at the contamination points in $\bm{t}=(t_{1},\cdots,t_{n})^{T}$, are, respectively, given by
@@
$IF(\bm{t},\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})=(1+\alpha)^{\frac{3}{2}}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{-1}\sum_{i=1}^{n}\nabla_{\bm{\beta}}\mu(\bm{x}_{i},\bm{\beta})(t_{i}-\mu(\bm{x}_{i},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i}-\mu(\bm{x}_{i},\bm{\beta}))^{2}},$ (18)
$IF(\bm{t},T_{\alpha}^{\sigma^{2}},\bm{F}_{\bm{\theta}})=\frac{2(1+\alpha)^{\frac{5}{2}}}{n(2+\alpha^{2})}\sum_{i=1}^{n}\left\{(t_{i}-\mu(\bm{x}_{i},\bm{\beta}))^{2}-\sigma^{2}\right\}e^{-\frac{\alpha}{2\sigma^{2}}(t_{i}-\mu(\bm{x}_{i},\bm{\beta}))^{2}}+\frac{2\alpha(1+\alpha)^{2}}{(2+\alpha^{2})}.$ (19)
@@
It is easy to observe that these IFs are also bounded in each argument $t_{i},\mbox{ for }i=1,2,\ldots,n$ for all $\alpha>0$, and unbounded at $\alpha=0$, as before.
@@
Let us now consider some IF-based summary measures of robustness for our NLR model. The gross error sensitivity of the minimum DPD functional $\bm{T^{\beta}}_{\alpha}$ of $\bm{\beta}$ at the vector of model distributions $\bm{F_{\theta}}$ with contamination only in the $i_{0}$-th observation is given by
@@
$\gamma_{i_{0}}^{u}(\bm{T}_{\alpha}^{\bm{\beta}},\bm{F}_{\bm{\theta}})$ $=\sup_{t}\left\{\lVert IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})\rVert\right\}$
@@
$=\left\{\begin{matrix}\frac{(1+\alpha)^{\frac{3}{2}}}{\sqrt{\alpha}}\sigma e^{-\frac{1}{2}}\lVert\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x}_{i_{0}},\bm{\beta})\rVert,&\mbox{if}&\alpha>0,\\ \infty,&\mbox{if}&\alpha=0.\end{matrix}\right$. (20)
But this measure does not satisfy the scale invariance property with respect to scale transformation of the components of $\bm{\beta}$. To overcome this problem, we consider a modified measure, namely the self-standardized sensitivity. For contamination only in the $i_{0}$-th observation, it is given by
$\gamma_{i_{0}}^{s}(\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})=\sup_{t}\left\{IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})^{T}(\bm{\Psi}_{pn}^{-1}\bm{\Omega}_{pn}\bm{\Psi}_{pn}^{-1})^{-1}IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})\right\}^{1/2}\\ =\left\{\begin{matrix}\frac{(1+2\alpha)^{\frac{3}{4}}}{\sqrt{n\alpha}}e^{-\frac{1}{2}}\left\{\nabla_{\bm{\beta}}\mu(\bm{x}_{i_{0}},\bm{\beta})^{T}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x}_{i_{0}},\bm{\beta})\right\}^{1/2},&\mbox{if}&\alpha>0,\\ \infty,&\mbox{if}&\alpha=0,\end{matrix}\right$. (21)
@@
where $\bm{\Psi}_{pn},\bm{\Omega}_{pn}$ are the $p\times p$ principal diagonal blocks of the matrices $\bm{\Psi}_{n},\bm{\Omega}_{n}$, respectively. Note that both the above sensitivity measures are decreasing functions in $\alpha\ (>0)$. So, it is evident that, with the presence of outliers, the robustness of the MDPDE of $\bm{\beta}$ increases with increasing values of $\alpha$. However, in Section 2.2 we have observed that the ARE of MDPDE of $\bm{\beta}$ (under pure data) decreases with increasing values of $\alpha$. Together, the tuning parameter $\alpha$ acts as a trade-off between the efficiency and robustness of the MDPDE under NLR models, consistent with the existing literature of DPD-based procedures.
@@
For contamination in all or some observations, the gross error sensitivity and the self-standardized sensitivity can similarly be defined as in (2.3) and (21), but replacing $IF_{i_{0}}(t,\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})$ by $IF(\bm{t},\bm{T}_{\alpha}^{\bm{\beta}},\bm{G})$ and taking supremum over all possible $t_{1}\ldots,t_{n}$. The resulting implications are again the same, and hence these details are omitted for brevity.